OpenStudy (anonymous):

a 10.0 kg ball falling at 10.0 m/s hits a mattress and comes to a complete stop in 0.5 s. What force does the mattress exert to stop the ball? A. 10 N B. 20 N C. 100 N D. 200 N @michele_laino :)

OpenStudy (michele_laino):

here we have to apply the theorem of the Impulse of a Force F. So we can write this: \[\Large F\Delta t = mv\] and then: \[\Large F = \frac{{mv}}{{\Delta t}} = \frac{{10 \times 10}}{{0.5}} = ...Newtons\]

OpenStudy (anonymous):

ok! so we get 200! so 200 N is our solutoin?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! :) thanks!

OpenStudy (michele_laino):

:)